# Monocular Lane Detection Based on Deep Learning: A Survey

### v1

@@ Abstract
Lane detection plays an important role in autonomous driving perception system. As deep learning algorithms gain popularity, monocular lane detection methods based on deep learning have demonstrated superior performance and emerged as a key research direction in autonomous driving perception. The core design of these algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global context supplementation, enhancing the detection of obscured lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. For a relatively fair comparison, in addition to comparing the performance of mainstream methods on different benchmarks, their inference speed is also investigated under a unified setting. Moreover, we present some extended works on lane detection, including multi-task perception, video lane detection, online high-definition (HD) map construction, and lane topology reasoning, to offer readers a comprehensive roadmap for the evolution of lane detection. Finally, we point out some potential future research directions in this field. We exhaustively collect the papers and codes of existing works at https://github.com/Core9724/Awesome-Lane-Detection and will keep tracing the research.
@@ I Introduction
Lane detection seeks to obtain the semantic and positional information of each lane line from the front view (FV) image captured by the onboard monocular camera. It is an indispensable part of the perception module in autonomous driving system, providing necessary prerequisites for subsequent decision-making and planning processes. Early lane detection methods depend on manually crafted operators for feature extraction [1, 2, 3, 4, 5, 6], which have limitations in accuracy and robustness when faced with complex scenes [7, 8, 9, 10, 11, 12, 13, 14, 15]. Later on, deep learning-based methods gradually dominate this field due to their strong feature representation ability and superior performance. Nowadays, deep learning-based monocular lane detection become a key research topic in autonomous driving perception ,attracting great attention from both academia and industry.
Figure 1: Two different technical routes to achieve complete lane detection. Different lane instances are visualized by different colors. The 2D lane detection results needs to be combined with camera parameters after which the conversion to 3D lanes is done with the help of IPM. 3D lane detection methods directly incorporate the camera parameters into the network and achieve end-to-end prediction of 3D lanes.
@@
Existing deep learning-based monocular lane detection methods can be divided into 2D lane detection and 3D lane detection methods. As shown in Figure 1, a complete lane detection process can be described as: given a FV image, the ultimate goal is to obtain 3D lane information in the ego vehicle coordinate system, i.e., the bird’s eye view (BEV) space. Because of the inherent perspective distortion in the camera imaging process, lanes that are parallel in the BEV plane intersect in the FV image, making it challenging to restore their true geometry. An effective solution involves building a 2D lane detection network [16, 17, 18, 19, 20] to obtain the 2D lanes from the FV image, after which the intrinsic and extrinsic parameters of the camera are combined to project these 2D lanes onto the ground through inverse perspective mapping (IPM) [21], thereby obtaining the final 3D lanes. With the rapid progress of deep learning-based methods, 2D lane detection achieve impressive results. However, IPM has an assumption that the ground is always flat, and does not consider the situations such as uphill, downhill, and rough road surfaces. In these cases, it lead to the inaccurate 3D lane results, even if the 2D lane detection in FV is correct. Therefore, researchers are gradually shifting their focus to designing 3D lane detection networks [22, 23, 24, 25, 26, 27], which directly predict 3D lanes using FV images as input.
@@
Similar to other perception tasks in autonomous driving, lane detection requires both precise lane positioning and real-time processing capabilities. Apart from the above, it also introduces distinct challenges. First, the network need to detect lanes which may be obscured due to nearby vehicles or severe weather conditions. Second, the network need to differentiate between various lane instances and present the lane detection results in a vectorized format, such as an ordered set of points or a curve equation. This is because downstream tasks require calculating the driving planning line, i.e. the centerline, based on the lanes on both sides of the vehicle. It is difficult to perform calculations without distinguishing between different lane instances or lacking a vectorized representation format.
@@
Based on the complete lane detection process and the unique task requirements, the core design of the lane detection algorithmic framework can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscured lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. A comprehensive survey on lane detection should systematically explore these four key perspectives while integrating insights from both 2D and 3D lane detection methods. This dual focus ensures a holistic understanding of the advancements in lane detection technology, bridging the conceptual design with practical applications.
@@
$\bullet$ Contribution. In order to provide readers better understand, we conduct a comprehensive investigation into the latest developments in monocular lane detection methods based on deep learning, starting from the core design of the lane detection algorithm framework. Compared to the related surveys [30, 31, 32], ours not only encompasses state-of-the-art 2D and 3D lane detection methods, but also summarizes at a more high level. The main contributions of this survey can be followed as:
@@
    This survey covers the core designs of four lane detection algorithms: task paradigm (Distinguishing different lane instances), lane modeling (representing lanes as network learnable parameters), global information supplementation (identifying obscured lanes), and eliminating perspective effects (obtaining available 3D lanes for downstream). In this way, each type of method can be summarized into a universal pipeline.
@@
    Moreover, some extended works are surveyed, including multi-task perception, video lane detection, online high-definition (HD) map construction and lane topology reasoning. They can be regarded as an upgrade of monocular lane detection in terms of task flow. Based on these introductions, readers can receive a roadmap for the development of lane detection research focus.
@@ II-A1 Datasets Commonly Used for 2D Lane Detection
$\bullet$ Tusimple. The TuSimple [35] dataset is collected with stable lighting conditions in highways, including different levels of occlusion, different types of lanes and different road conditions. It only consists of 6,408 images, which are split into 3,268 training, 358 validation, and 2,782 test images. For each image, lanes are annotated by the 2D coordinates of sampling points with a uniform height interval of 10 pixels. Each annotated image has a size of 1280$\times$720 pixels.
$\bullet$ CULane. CULane [36] is a large-scale 2D lane detection dataset with 88,880 training images and 34,680 testing images. In addition to different weather conditions and light levels, there are eight challenging lane detection scenarios, such as traffic congestion, shadow occlusion, missing lanes and lane curves. All the images have 1640$\times$590 pixels.
$\bullet$ LLAMAS. The annotations of LLAMAS [39] are automatically genetrated from HD maps. This dataset contains over 100k images from about 350 km recorded drives. In contrast to other datasets, LLAMAS presents a small and variable number of pixels marking each lane, reflecting real-world conditions more accurately. The resolution of all images is 1280$\times$717 pixels.
$\bullet$ CurveLanes. CurveLanes [40] contains 100K, 20K, and 30K images for training, validation, and testing, respectively. It features an abundance of curved lanes and difficult scenarios such as S-curves, fork lines, night time conditions, and multi-lane configurations. In comparison to existing datasets like the first three, each image within CurveLanes may encompass a greater number of lanes and has higher resolution.
@@ II-A2 Datasets Commonly Used for 3D Lane Detection
$\bullet$ Apollo 3DLane. The Apollo 3DLane dataset [23] is generated using game engine, including 10,500 discrete frames of monocular RGB images and their corresponding 3D lanes ground truth, which is split into three scenes: balanced, rarely observed, and visual variation scenes. Each scene contains independent training sets and test sets.
$\bullet$ ONCE-3DLanes. ONCE-3DLanes [44] is a large-scale real-world 3D lane detection dataset, which is constructed based on the ONCE dataset [47]. It contains 211K images comprising diverse weather conditions (sunny, cloudy, rainy) and varied geographical locations (urban centers, suburban areas, highways, bridges and tunnels). Only internal camera are provided in ONCE-3DLanes.
$\bullet$ OpenLane. OpenLane [24] is an another large-scale but more comprehensive benchmark for real-world 3D lane detection based on Waymo Open Dataset [48]. The dataset includes 200K images captured in a variety of weather, terrain, and brightness conditions. In OpenLane, the lane annotation not only contain the 3D position of a lane, but also several attribute and tracking id. The intrinsics and extrinsics of the camera are provided for each frame, and category and scene labels (e.g., weather and location) are also provided, providing a realistic and diverse set of challenges for 3D lane detection algorithms.
@@ II-B Evaluation Metrics
As 2D lane detection benchmark representives, Tusimple [35] and CULane [36] adopts two different ways to determine TP. Tusimple [35] focuses on point-by-point evaluation. The predicted point is considered correct if the horizontal distance from the true value point is less than 20 pixels when the longitudinal coordinates are the same. Furthermore, the line prediction is viewed as TP when it contains no less than 85% of the true value points. In contrast, CULane [36] emphasizes line-by-line evaluation, treating each lane as a mask of several pixels wide, and calculating the intersection (IoU) between the predicted lane and the annotated lane. The prediction with IoU larger than 75% is viewed as TP.
@@ III Methods of 2D Lane Detection
This section reviews the existing 2D lane detection methods. We first explain the ground for classifying existing methods in Section III-A, and then discuss the existing methods accordingly in Section III-B and Section III-C. Lastly, the process of converting 2D lanes to 3D lanes using IPM and its deficiencies are reviewd in Section III-D.
@@ III-A Classification Framework
Previous summaries [30, 31] primarily focus on the design of network structures and loss functions. It is noteworthy that instance-level discrimination and vectorized result representation, as unique requirements for the lane detection task and necessary prerequisites for guiding downstream applications, have been overlooked. By contrast, our classification of 2D lane detection methods are primarily according to the above two aspects.
As shown in Figure 3, first, for lane instance-level discrimination, 2D lane detection methods can be divided into two types of paradigms based on the number of stages required to complete the task: (a) Segmentation-based methods (two-stage), which complete lane recognition and instance discrimination in a certain order. Figure 4 summarize the general pipeline for instance discrimination in such methods. (b) Object detection-based methods (one-stage), which can directly perform instance discrimination and localization concurrently. This advantage arises from the general pipeline of object detection algorithms, which execute both classification and regression tasks on a set of candidate proposals in parallel.
Second, vectorized result representation requires algorithms to consider how to model lanes as a set of values for neural network learning, i.e., lane modeling. In terms of lane modeling, the segmentation-based methods can be further divided into mask-based modeling, grids-based modeling and keypoints-based modeling. For object detection-based methods, adopting bounding boxes (bbox) to model narrow and long lanes is often not reasonable. This is because the bbox generated by object detection methods are mutually occluded, and there may be multiple instances of the same type of lane in the bbox. To align with the general object detection paradigm, these methods design unique ”bbox” to model lanes, including line anchor-based modeling and curve-based modeling. The details of each lane modeling are described in Figure 5.
Figure 3: The genaral piplines for 2D lane detection: (a) Segmentation-based methods, which leverage mask, grids or keypoints to model lanes. Each method only performs instance discrimination once. (b) Object detection-based methods, which leverage line anchor or parameter curve to model lanes and reach instance discrimination and localization in parallel.
Furthermore, most existing 2D lane detection datasets provide complete annotations for lanes, even in cases where the lanes are severely occluded by vehicles or affected by extreme weather conditions. This prompts many algorithms to meticulously design special structures within their networks to address this challenge, regardless of which classification system they belong to. Therefore, while reviewing these methods, the measures which they have taken to tackle it are also be elaborated. We elobrately compare the representive 2D lane detection works according to the above classification criteria in Table II.
@@ III-B1 Mask-based modeling
SCNN [36] proposes a top-down manner. The approach is treating each lane in the image as a separate category and performing multi-category semantic segmentation. It incorporates a parallel classification branch to predict the existence of lanes at each position, subsequently combining the classification and segmentation results during post-processing to obtain the final lane. The subsequent work [58, 56, 16, 61] followed this method. However, these methods facilitates instance differentiation but introduces certain limitations: it requires defining a maximum number of lanes in advance to determine the number of possible instances. Additionally, the correspondence between lanes and classes is established by annotations. When vehicles switch between lanes, this predefined labeling may lead to ambiguity.
Figure 4: The ways for discriminating different lane instances in segmentation-based methods: (a) Bottom-up faction represented by [65, 66, 67, 18]. (b) Top-down faction represented by [36, 56, 16, 68, 18], which predefine the maximum number of lanes and treat each lane as a category. (c) Top-down faction represented by [17, 69], which predict dynamic kernels to generate instances. Among them, grids-based modeling is only applicable to the top-down approach.
To slove the above problem, some studies adopt a bottom-up approach for a more flexible instance segmentation, e.g, cluster the binary segmentation results of lanes/backgrounds. VPGNet [34] clusters lanes using a modified density-based clustering method. LaneNet [65] utilizes instance embedding to cluster the results of semantic segmentation, achieving lane instance segmentation. This method offers high clustering accuracy but is time-consuming, which limits its applicability for real-time vehicle experiments. FastDraw [70] constructs a learnable decoder that not only segments lanes but also identifies pixels belonging to the same lane. To address the inefficiency of pixel-embedding clustering, LaneAF [66] introduces an affinity vector field to associate pixels belonging to the same lane. Although these methods improve instance discrimination, the algorithm execution efficiency remains suboptimal due to the high complexity of bottom-up clustering and the low efficiency of mask-based modeling in classifying all pixels.
The lane mask obtained from network segmentation usually contains a large number of irrelevant areas. In order to be used for self vehicle motion prediction and planning, it is necessary to further denoise the mask to obtain vectorized results. Usually, for each lane mask, the highest response is sampled sequentially at equidistant heights, and then curve fitting is performed.

### v2

@@ Abstract
Lane detection plays an important role in autonomous driving perception systems. As deep learning algorithms gain popularity, monocular lane detection methods based on deep learning have demonstrated superior performance and emerged as a key research direction in autonomous driving perception. The core design of these algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global context supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. For a relatively fair comparison, in addition to comparing the performance of mainstream methods on different benchmarks, their inference speed is also investigated under a unified setting. Moreover, we present some extended works on lane detection, including multi-task perception, video lane detection, online high-definition map construction, and lane topology reasoning, to offer readers a comprehensive roadmap for the evolution of lane detection. Finally, we point out some potential future research directions in this field. We exhaustively collect the papers and codes of existing works at https://github.com/Core9724/Awesome-Lane-Detection and will keep tracing the research.
@@ I Introduction
Lane detection seeks to obtain the semantic and positional information of each lane line from the front view (FV) image captured by the onboard monocular camera. It is an indispensable part of the perception module in autonomous driving system, providing necessary prerequisites for subsequent decision-making and planning processes. Early lane detection methods depend on manually crafted operators for feature extraction [1, 2, 3, 4, 5, 6], which have limitations in accuracy and robustness when faced with complex scenes [7, 8, 9, 10, 11, 12, 13, 14, 15]. Later on, deep learning-based methods gradually dominate this field due to their strong feature representation ability and superior performance. Nowadays, deep learning-based monocular lane detection becomes a key research topic in autonomous driving perception, attracting great attention from both academia and industry.
Figure 1: Two different technical routes to achieve complete lane detection. Different lane instances are visualized by different colors. The 2D lane detection results need to be combined with camera parameters after which the conversion to 3D lanes is done with the help of IPM. 3D lane detection methods directly incorporate the camera parameters into the network and achieve end-to-end prediction of 3D lanes.
@@
Existing deep learning-based monocular lane detection methods can be divided into 2D lane detection and 3D lane detection methods. As shown in Figure 1, a complete lane detection process can be described as: given a FV image, the ultimate goal is to obtain 3D lane information in the ego vehicle coordinate system, i.e., the bird’s eye view (BEV) space. Because of the inherent perspective distortion in the camera imaging process, parallel lanes in the BEV plane intersect in the FV image, making it challenging to restore their true geometry. An effective solution involves building a 2D lane detection network [16, 17, 18, 19, 20] to obtain the 2D lanes from the FV image, after which the intrinsic and extrinsic parameters of the camera are combined to project these 2D lanes onto the ground through inverse perspective mapping (IPM) [21], thereby obtaining the final 3D lanes. With the rapid progress of deep learning-based methods, 2D lane detection achieve impressive results. However, IPM has an assumption that the ground is always flat, and does not consider the situations such as uphill, downhill, and rough road surfaces. In these cases, it leads to inaccurate 3D lane results, even if the 2D lane detection in FV is correct. Therefore, researchers are gradually shifting their focus to designing 3D lane detection networks [22, 23, 24, 25, 26, 27], which directly predict 3D lanes using FV images as input.
@@
For stable driving of vehicles in complex road environments, lane detection requires both precise localization and real-time processing capabilities. Apart from the above, the lanes which may be obscure due to nearby vehicles or severe weather conditions need to be detected. In order to better connect downstream tasks such as planning and control, the network also needs to differentiate between various lane instances and present the lane detection results in a vectorized format, such as an ordered set of points or a curve equation. This is because downstream requires calculating the driving planning line, i.e. the centerline, based on the lanes on both sides of the vehicle. It is difficult to perform calculations without distinguishing between different lane instances or lacking a vectorized representation format.
@@
Based on the complete lane detection process and the unique task requirements, the core design of the lane detection algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. A comprehensive survey on lane detection should systematically explore these four key perspectives while integrating insights from both 2D and 3D lane detection methods. This dual focus ensures a holistic understanding of the advancements in lane detection technology, bridging the conceptual design with practical applications.
@@
$\bullet$ Contribution. To provide readers better understanding, we conduct a comprehensive investigation into the latest developments in monocular lane detection methods based on deep learning, starting from the core design of the lane detection algorithm frameworks. Compared to the related surveys [30, 31, 32], ours not only encompasses state-of-the-art 2D and 3D lane detection methods but also summarizes at a higher level. The main contributions of this survey can be followed as:
@@
    This survey covers the core designs of four-lane detection algorithms: task paradigm (Distinguishing different lane instances), lane modeling (representing lanes as network learnable parameters), global information supplementation (identifying obscure lanes), and eliminating perspective effects (obtaining available 3D lanes for downstream). In this way, each type of method can be summarized into a universal pipeline.
@@
    Moreover, some extended works are surveyed, including multi-task perception, video lane detection, online high-definition (HD) map construction, and lane topology reasoning. They can be regarded as an upgrade of monocular lane detection in terms of task flow. Based on these introductions, readers can receive a roadmap for the development of lane detection research focus.
@@ II-A1 Datasets Commonly Used for 2D Lane Detection
$\bullet$ Tusimple. The TuSimple [35] dataset is collected with stable lighting conditions in highways, including different levels of occlusion, different types of lanes, and different road conditions. It only consists of 6,408 images, which are split into 3,268 training, 358 validation, and 2,782 test images. For each image, lanes are annotated by the 2D coordinates of sampling points with a uniform height interval of 10 pixels. Each annotated image has a size of 1280$\times$720 pixels.
$\bullet$ CULane. CULane [36] is a large-scale 2D lane detection dataset with 88,880 training images and 34,680 testing images. In addition to different weather conditions and light levels, there are eight challenging lane detection scenarios, such as traffic congestion, shadow occlusion, missing lanes, and lane curves. All the images have 1640$\times$590 pixels.
$\bullet$ LLAMAS. The annotations of LLAMAS [39] are automatically generated from HD maps. This dataset contains over 100k images from about 350 km of recorded drives. In contrast to other datasets, LLAMAS presents a small and variable number of pixels marking each lane, reflecting real-world conditions more accurately. The resolution of all images is 1280$\times$717 pixels.
$\bullet$ CurveLanes. CurveLanes [40] contains 100K, 20K, and 30K images for training, validation, and testing, respectively. It features an abundance of curved lanes and difficult scenarios such as S-curves, fork lines, nighttime conditions, and multi-lane configurations. In comparison to existing datasets like the first three, each image within CurveLanes encompasses a greater number of lanes and has a higher resolution.
@@ II-A2 Datasets Commonly Used for 3D Lane Detection
$\bullet$ Apollo 3DLane. The Apollo 3DLane dataset [23] is generated using the game engine, including 10,500 discrete frames of monocular RGB images and their corresponding 3D lanes ground truth, which is split into three scenes: balanced, rarely observed, and visual variation scenes. Each scene contains independent training sets and test sets.
$\bullet$ ONCE-3DLanes. ONCE-3DLanes [44] is a large-scale real-world 3D lane detection dataset, which is constructed based on the ONCE dataset [47]. It contains 211K images comprising diverse weather conditions (sunny, cloudy, rainy) and varied geographical locations (urban centers, suburban areas, highways, bridges, and tunnels). Only intrinsics of the camera are provided in ONCE-3DLanes.
$\bullet$ OpenLane. OpenLane [24] is another large-scale but more comprehensive benchmark for real-world 3D lane detection based on Waymo Open Dataset [48]. The dataset includes 200K images captured in a variety of weather, terrain, and brightness conditions. In OpenLane, the lane annotation not only contains the 3D position of a lane but also several attributes and tracking id. The intrinsics and extrinsics of the camera are provided for each frame, and category and scene labels (e.g., weather and location) are also provided, providing a realistic and diverse set of challenges for 3D lane detection algorithms.
@@ II-B Evaluation Metrics
As 2D lane detection benchmark representatives, Tusimple [35] and CULane [36] adopt two different ways to determine TP. Tusimple [35] focuses on point-by-point evaluation. The predicted point is considered correct if the horizontal distance from the true value point is less than 20 pixels when the longitudinal coordinates are the same. Furthermore, the line prediction is viewed as TP when it contains no less than 85% of the true value points. In contrast, CULane [36] emphasizes line-by-line evaluation, treating each lane as a mask of several pixels wide, and calculating the intersection (IoU) between the predicted lane and the annotated lane. The prediction with IoU larger than 75% is viewed as TP.
@@ III Methods of 2D Lane Detection
This section reviews the existing 2D lane detection methods. We first explain the ground for classifying existing methods in Section III-A, and then discuss the existing methods accordingly in Section III-B and Section III-C. Lastly, the process of converting 2D lanes to 3D lanes using IPM and its deficiencies are reviewed in Section III-D.
@@ III-A Classification Framework
Previous summaries [30, 31] primarily focus on the design of network structures and loss functions. It is noteworthy that instance-level discrimination and vectorized result representation, as unique requirements for the lane detection task and necessary prerequisites for guiding downstream applications, have been overlooked. By contrast, our classification of 2D lane detection methods is primarily according to the above two aspects.
As shown in Figure 3, first, for lane instance-level discrimination, 2D lane detection methods can be divided into two types of paradigms based on the number of stages required to complete the task: (a) Segmentation-based methods (two-stage), which complete lane recognition and instance discrimination in a certain order. Figure 4 summarizes the general pipeline for instance discrimination in such methods. (b) Object detection-based methods (one-stage), which can directly perform instance discrimination and localization concurrently. This advantage arises from the general pipeline of object detection algorithms, which execute both classification and regression tasks on a set of candidate proposals in parallel.
Second, vectorized result representation requires algorithms to consider how to model lanes as a set of values for neural network learning, i.e., lane modeling. In terms of lane modeling, the segmentation-based methods can be further divided into mask-based modeling, grids-based modeling, and keypoints-based modeling. For object detection-based methods, adopting a bounding box (bbox) to model a narrow and long lane is often not reasonable. This is because the bbox generated by object detection methods are mutually occluded, and there may be multiple instances of the same type of lane in the bbox. To align with the general object detection paradigm, these methods design unique ”bbox” to model lanes, including line anchor-based modeling and curve-based modeling. The details of each lane modeling are described in Figure 5.
Figure 3: The general pipelines for 2D lane detection: (a) Segmentation-based methods, which leverage mask, grids, or keypoints to model lanes. Each method only performs instance discrimination once. (b) Object detection-based methods, which leverage line anchor or parameter curve to model lanes and reach instance discrimination and localization in parall
Furthermore, most existing 2D lane detection datasets provide complete annotations for lanes, even in cases where the lanes are severely occluded by vehicles or affected by extreme weather conditions. This prompts many algorithms to meticulously design special structures within their networks to address this challenge, regardless of which classification system they belong to. Therefore, while reviewing these methods, the measures which they have taken to tackle it are also elaborated. We elaborately compare the representative 2D lane detection works according to the above classification criteria in Table II.
@@ III-B1 Mask-based modeling
SCNN [36] proposes a top-down manner. The approach is treating each lane in the image as a separate category and performing multi-category semantic segmentation. It incorporates a parallel classification branch to predict the existence of lanes at each position, subsequently combining the classification and segmentation results during post-processing to obtain the final lane. The subsequent work [58, 56, 16, 61] followed this method. However, these methods facilitate instance differentiation but introduce certain limitations: it requires defining a maximum number of lanes in advance to determine the number of possible instances. Additionally, the correspondence between lanes and classes is established by annotations. When vehicles switch between lanes, this predefined labeling may lead to ambiguity.
Figure 4: The ways for discriminating different lane instances in segmentation-based methods: (a) Bottom-up faction represented by [65, 66, 67, 18]. (b) Top-down faction represented by [36, 56, 16, 68, 18], which predefines the maximum number of lanes and treats each lane as a category. (c) Top-down faction represented by [17, 69], which predicts dynamic kernels to generate instances. Among them, grids-based modeling is only applicable to the top-down approach.
To solve the above problem, some studies adopt a bottom-up approach for a more flexible instance segmentation, e.g., cluster the binary segmentation results of lanes/backgrounds. VPGNet [34] clusters lanes using a modified density-based clustering method. LaneNet [65] utilizes instance embedding to cluster the results of semantic segmentation, achieving lane instance segmentation. This method offers high clustering accuracy but is time-consuming, which limits its applicability for real-time vehicle experiments. FastDraw [70] constructs a learnable decoder that not only segments lanes but also identifies pixels belonging to the same lane. To address the inefficiency of pixel-embedding clustering, LaneAF [66] introduces an affinity vector field to associate pixels belonging to the same lane. Although these methods improve instance discrimination, the algorithm execution efficiency remains suboptimal due to the high complexity of bottom-up clustering and the low efficiency of mask-based modeling in classifying all pixels.
The lane mask obtained from network segmentation usually contains a large number of irrelevant areas. In order to be used for ego-vehicle motion prediction and planning, it is necessary to further denoise the mask to obtain vectorized results. Usually, for each lane mask, the highest response is sampled sequentially at equidistant heights, and then curve fitting is performed.